OpenStudy (anonymous):

Solve the formula for thee indicated variable : Formula for the flow of current in an electrical circuit, V+IH, for I

OpenStudy (anonymous):

you need an equal sign somewheres to have an equation to solve you wote \[V+IH\] which is an expression

OpenStudy (anonymous):

was it supposed to be \[V=IH\]?

OpenStudy (anonymous):

yes sorry misstyped it

OpenStudy (anonymous):

ok it is like if \(12=4\times 3\) then \(4=\frac{12}{3}\)

OpenStudy (anonymous):

\[V=IH, I=\frac{V}{H}\]

OpenStudy (anonymous):

howd u get that

OpenStudy (anonymous):

tried to explain with the first example i did with numbers do you see it?

OpenStudy (anonymous):

\[75=3\times 25\\ 3=\frac{75}{25}\] they say the same thing

OpenStudy (anonymous):

oh ok so the V goes on top always and the H on the bottom,

OpenStudy (anonymous):

your math teacher would say "divide both sides by \(H\) " but really you just write it

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw